OpenStudy (anonymous):

A 1300 g mass is on a horizontal surface with mu-k = 0.470, and is in contact with a massless spring with a force constant of 535 N/m which is compressed. When the spring is released, it does 2.68 J of work on the mass while returning to its equilibrium position. Calculate the distance the spring was compressed.

OpenStudy (shamim):

Required formula is W=0.5*k*x^2 Given W=2.68J k=535N/m x=?